Determine the measure of angle B of triangle ACB, given the following information: A=69degrees, a=25ft, b=16ft
Answer by throwabunny(3)

Reckon they're looking for law of sin there.
Which is.
a/sin A = b/sin B =c/sin C
So just plug in stuff.
25/sin 69 degrees = 16/sin x degrees
and solve.
I got something along the lines of 36.69 degrees.